massachusetts institute of technology

Department of Electrical Engineering and Computer Science

Graduate Office

Area II Minor Requirement (Old Program)

One of the requirements or doctoral study in the Department of Electrical Engineering and Computer Science is the completion of a minor program. The object of the minor program is to broaden the student’s grasp of a field outside the field of major professional interest; often the student is exposed to methodologies and ways of thinking that are different from those encountered in the major field. For an experimentalist, for example, the abstract concepts of mathematics form such a different methodology; for a theoretician, the new methodology might be the practical considerations of engineering.

The minor program should be a cohesive program consisting of at least three subjects that are closely related in content with a minimum of 27 units. For example, three subjects in a single area like Topology, Accounting, Solid State Physics, or Neurophysiology are acceptable selections. It should be pointed out that a collection of subjects offered by the same academic department should not automatically be considered to be a cohesive package. On the other hand, a selection of subjects from different departments may well form an acceptable program.

The minor program should be a study program outside of the major professional area of study. As an example, Computer Science subjects are normally not appropriate as part of a minor program for students in Area II, but advanced study in areas close to Computer Science may b acceptable as a minor program. The decision in such case should be based on the criterion of whether the studies in question expose the student to a methodology and a way of thinking that are different from those found in the major study program.

The minor program should involve study beyond the introductory level although. when appropriate, the program may contain one subject at an undergraduate or introductory graduate level. It should be noted that certain subjects with graduate ’H’ credit are considered to be introductory. The Graduate Office should be consulted if there are questions concerning the acceptability of certain subjects. Subjects in the series 6.981 and 6.991 series are not acceptable.

It is a departmental requirement that at least one of the subjects in a minor program must be taken on campus at MIT. No subjects with a grade of C or lower will be accepted in a minor program.

Minor study may start as early in the graduate program as the student wishes. However, it is strongly advised that the suitability of the proposed program be discussed with the Graduate Counselor before beginning minor studies. The program should be agreed to by the Counselor before it is commenced. The program must also have the approval of the Area Committee acting for the Department Committee on Graduate Students. Hence, the plan for the minor should be included in the overall doctoral plan submitted to the Graduate Office through the Counselor on Form 3910. The field of the minor as well as the field of the major must be named in the plan submitted to the Graduate Office so that the distinction between them is clear. Revisions in the student’s plan for major study may require that a new minor program be formulated and approved.

A student’s minor program should be formulated when applying for the Oral Qualifying Examination for doctoral study. This insures that the minor program will be included in the plan for graduate study when it is submitted to the Electrical Engineering and Computer Science Graduate Office.

Minor programs for students in each Area will be evaluated and certified by that Area Committee at the time the student is judged “Qualified” for the doctoral program on completion of the Oral Qualifying Examination.